Solve the absolute value equation.

| x + |3x - 2| | = 2


Here is my set-up:



| x + |3x - 2| | = 2 or | x + |3x - 2| | = -2


Is this the correct first step?


If so, what is the next step?

Answer by ikleyn(53937)

.

I will answer exactly as you ask in your post.


Your first step setup is incorrect.

The correct first step should be as follows

    x + |3x-2| = 2  or  x + |3x-2| = -2.

Pay attention on how I take off the absolute value signs:
I take them off (very extreme absolute value signs), but I write in the right side +2 or -2.


-----------

As the next step,  you should consider two cases:

        (a)   3x-2 >= 0,     and

        (b)   3x-2 < 0.


***********************************************

                I just answered your questions.

***********************************************


In this web-site www.algebra.com,  there is a group of lessons,  where I show/explain how to treat similar equations.

These lessons are
    - HOW TO solve equations containing Linear Terms under the Absolute Value sign. Lesson 1
    - HOW TO solve equations containing Linear Terms under the Absolute Value sign. Lesson 2
    - HOW TO solve equations containing Linear Terms under the Absolute Value sign. Lesson 3

It is the material for you to learn from.